OpenStudy (anonymous):

please help... if 3sinθ + 4cosθ = 5, 3cosθ-4sinθ = ?

OpenStudy (anonymous):

comt to twiddla http://www.twiddla.com/583597

OpenStudy (anonymous):

wt is dat?

OpenStudy (anonymous):

click on it

OpenStudy (anonymous):

it is like a board so that i can explain it to u easily

OpenStudy (anonymous):

k..

OpenStudy (anonymous):

are u here........................

OpenStudy (anonymous):

Light has finally dawned....

OpenStudy (anonymous):

This is just a 3,4,5 triangle with sin =3/5 and cos = 4/5. Zero.

OpenStudy (anonymous):

how did u do that estudier? can u pls tell me

OpenStudy (anonymous):

First I did all the usual things, substitution, cross multiplication, trig things and I always had a theta in the answer. Then I thought about triangles and I saw almost straight away that a 3,4,5 Pythagorean triangle was the way to the answer (I should have thought of it sooner because there is a 3,4,5 in the question:-)

OpenStudy (anonymous):

so how will it be a part of the triangle? i can't find a connection..pls help

OpenStudy (anonymous):

What is sin theta and cos theta in a 3,4,5 triangle?

OpenStudy (anonymous):

sin theta= 3/5 cos theta= 4/5

OpenStudy (anonymous):

thank u..:) how did u guess it from the question?

OpenStudy (anonymous):

I explained in my post above.

OpenStudy (anonymous):

First I did all the usual things, substitution, cross multiplication, trig things and I always had a theta in the answer. Then I thought about triangles and I saw almost straight away that a 3,4,5 Pythagorean triangle was the way to the answer (I should have thought of it sooner because there is a 3,4,5 in the question:-)

OpenStudy (anonymous):

thanks a lot.... :)

OpenStudy (anonymous):

u r welcome.